### v2

# Implications of recent Higgs data for the light Higgs bosons in the GNMSSM

Abstract
This study systematically investigates the parameter space of the GNMSSM featuring light Higgs bosons, with a focus on the exotic decay of the 125 GeV Higgs boson ($h$) into pairs of the lightest CP-odd ($a_{1}$) or CP-even ($h_{1}$) Higgs bosons under constraints from recent Higgs data. We perform an initial scan of the parameter space under basic constraints from HiggsSignals-2.6.2 and HiggsBounds-5.10.2, followed by the application of comprehensive constraints using the HiggsTools framework. Comparative analysis reveals that HiggsTools imposes significantly more stringent constraints than its legacy components alone. Furthermore, its exclusion capability substantially surpasses that of ATLAS direct searches targeting individual decay channel.
By analyzing the distinct constraining effects of the subpackages HiggsSignals and HiggsBounds within HiggsTools, we find HiggsBounds constraints to be dominant across both the $h_{1}$ and $h_{2}$ scenarios. Nearly all parameter points excluded by HiggsSignals or direct LHC decay channel analyses are also excluded by HiggsBounds, attributable to the direct constraints HiggsBounds provides for non-SM Higgs searches, in contrast to the indirect constraints from HiggsSignals. In the $h_{2}$ scenario, HiggsSignals can exclude points with suppressed $Br(h\to a_{1}a_{1}\to\tau\tau bb)$ due to the kinematically accessible decay $h_{2}\to h_{1}h_{1}$, which exhibits large branching fractions and modifies SM-like Higgs properties. Our analysis shows that experimental data require the observed $h$ boson to be predominantly SM-like in both scenarios. Specifically, in the $h_{2}$ scenario, the lighter singlet-dominated state $h_{s}$ must possess high singlet purity and negligible SM-like components.
1 Introduction
The discovery of the 125 GeV Higgs boson at the Large Hadron Collider (LHC) [1, 2] in 2012 marked a monumental milestone in particle physics, confirming the mechanism of electroweak symmetry breaking (EWSB) as described by the Standard Model (SM). However, despite its tremendous success, the SM leaves several fundamental questions unresolved, such as neutrino masses, the hierarchy problem, and the dark matter problem. These unresolved issues have motivated extensive exploration of physics beyond the Standard Model (BSM), where extended Higgs sectors and new symmetries often predict additional scalar states.
Supersymmetry (SUSY) stands out as a compelling BSM candidate, offering solutions like addressing the hierarchy problem and enabling gauge coupling unification. The Minimum Supersymmetric Standard Model (MSSM) [3, 4, 5, 6], as the most economical SUSY realization, has been extensively studied. Yet, it faces significant limitations from dark matter constraints and struggles with issues like the $\mu$ problem and the little hierarchy problem. As a natural and minimal extension, the Next-to-Minimal Supersymmetric Standard Model (NMSSM) [7, 8, 9] introduces a gauge-singlet Higgs superfield. Its scalar component acquires a non-zero vacuum expectation value (vev), dynamically generating an effective $\mu$-term. Crucially, mixing between the singlet and doublet superfields produces a new CP-even Higgs ($h_{s}$) that mixes with MSSM-like states. This mixing can suppress the lightest Higgs mass while allowing the next-to-lightest Higgs to match the observed 125 GeV signal [10, 11, 12, 13, 14, 15, 16]. This scenario alleviates fine-tuning and offers rich phenomenology, including Higgs-to-Higgs decays [17, 18, 19, 20, 21, 22, 23, 24] or exotic decays involving neutralinos [25].
However, the predictive power of the $Z_{3}$-invariant NMSSM is constrained by its superpotential structure, motivating generalizations like the General NMSSM (GNMSSM) [26, 27, 28, 29, 30, 31]. By adding extra singlet couplings or $Z_{3}$-symmetry-breaking terms, the GNMSSM enhances flexibility in the Higgs mass matrix, enabling lighter CP-even and CP-odd states consistent with LHC data. These light Higgs bosons could evade direct detection due to suppressed SM couplings but might appear through exotic decays of the 125 GeV Higgs $h$, such as $h\to h_{s}h_{s}/a_{1}a_{1}$ (where $a_{1}$ is the lightest CP-odd Higgs). Such rare decays transform the 125 GeV Higgs into a unique portal for discovering light hidden sector particles.
Both the ATLAS and CMS collaborations at the LHC have pursued extensive searches for light Higgs bosons through exotic decays of the 125 GeV Higgs boson [32], targeting final states like $4b$ [33], $\mu\mu bb$ [34, 35, 36], $\tau\tau bb$ [37, 38], $\gamma\gamma\tau\tau$ [39], and $4\mu$ [40, 41, 42, 43, 44]. Despite stringent limits from these searches, no significant excess has been observed, imposing strong constraints on the GNMSSM parameter space. This work primarily focuses on the promising yet challenging $\tau\tau bb$ final state [38]. Furthermore, we incorporate complementary constraints on the parameter space using the package HiggsTools [45], which combines and extends the functionalities of HiggsBounds and HiggsSignals. HiggsBounds tests BSM model compatibility against exclusion limits from new scalar searches, while HiggsSignals assesses predictions against experimental measurements of the 125 GeV Higgs boson production rates.
This study presents a comprehensive investigation of the 125 GeV Higgs boson’s exotic decays into pairs of lighter CP-even or CP-odd Higgs bosons within the GNMSSM framework. By systematically incorporating the stringent constraints from LHC direct Higgs searches and HiggsTools, we explore the viable parameter space capable of accommodating such signatures in the GNMSSM.
This paper is structured as follows. Section 2 introduces the GNMSSM framework, detailing its Higgs sector and the key trilinear Higgs couplings essential for our analysis. In Section 3, we outline our scanning strategy and present the corresponding numerical results. Finally, Section 4 summarizes our conclusions.
2 Theoretical preliminaries
2.1 Basics of the GNMSSM
The GNMSSM extends the MSSM by introducing a gauge-singlet superfield $\hat{S}$, carrying neither baryon nor lepton number. Consequently, the Higgs sector comprises not only the standard $SU(2)_{L}$ doublets $\hat{H}_{u}=(\hat{H}^{+}_{u},\hat{H}^{0}_{u})$ and $\hat{H}_{d}=(\hat{H}^{0}_{d},\hat{H}^{-}_{d})$, but also the singlet $\hat{S}$. Its gauge-invariant superpotential is given by  [8]:
$W_{\rm GNMSSM}=W_{\text{Yukawa}}+\lambda\hat{S}\hat{H_{u}}\cdot\hat{H_{d}}+\frac{\kappa}{3}\hat{S}^{3}+\mu\hat{H_{u}}\cdot\hat{H_{d}}+\frac{1}{2}\mu^{\prime}\hat{S}^{2}+\xi\hat{S},$ (2.1)
where $W_{\rm Yukawa}$ contains the MSSM quark and lepton Yukawa interactions. The dimensionless couplings $\lambda$ and $\kappa$ govern Higgs interactions, analogous to the $Z_{3}$-NMSSM. Crucially, the GNMSSM incorporates explicit $Z_{3}$-symmetry-breaking via the parameters $\mu$, $\mu^{\prime}$ and $\xi$. These terms resolve the tadpole problem [46, 8] and the cosmological domain-wall problem [47, 48, 49] inherent in the $Z_{3}$-invariant NMSSM. The parameter $\xi$ can be consistently eliminated by singlet field redefinition  [50], allowing us to set $\xi=0$ without loss of generality. Furthermore, $\mu$ and $\mu^{\prime}$ naturally arise from the spontaneous breaking of discrete $R$-symmetries ($\mathbb{Z}^{R}_{4}$ or $\mathbb{Z}^{R}_{8}$) at high scales [47, 51, 52, 50, 53]. These explicit $Z_{3}$-breaking terms significantly modify the neutral Higgs mass spectrum, generating phenomenology markedly richer than both the $Z_{3}$-NMSSM and the MSSM.
2.2 The Higgs Sector of GNMSSM
The soft SUSY-breaking terms in the Higgs sector of the GNMSSM can be expressed as:
$-\mathcal{L}_{soft}=$ $\Bigg{[}\lambda A_{\lambda}SH_{u}\cdot H_{d}+\frac{1}{3}\kappa A_{\kappa}S^{3}+m_{3}^{2}H_{u}\cdot H_{d}+\frac{1}{2}{m_{S}^{\prime}}^{2}S^{2}+\xi^{\prime}S+h.c.\Bigg{]}$ (2.2)
$+m^{2}_{H_{u}}|H_{u}|^{2}+m^{2}_{H_{d}}|H_{d}|^{2}+m^{2}_{S}|S|^{2},$
with $H_{u},H_{d},$ and $S$ denoting the scalar components of the Higgs superfields, and $m^{2}_{H_{u}},m^{2}_{H_{d}},$ and $m^{2}_{S}$ being their supersymmetry-breaking masses. Following the electroweak symmetry breaking, the neutral components of the Higgs fields develop non-vanishing vevs:
$\langle H^{0}_{u}\rangle=\frac{v_{u}}{\sqrt{2}},\quad\langle H^{0}_{d}\rangle=\frac{v_{d}}{\sqrt{2}},\quad\langle S\rangle=\frac{v_{s}}{\sqrt{2}},$ (2.3)
with $v=\sqrt{v_{u}^{2}+v_{d}^{2}}\simeq 246\,\text{GeV}$. The Higgs sector is then characterized by eleven independent physical parameters:
$\tan\beta\equiv\frac{v_{u}}{v_{d}},\,\lambda,\,\kappa,\,v_{s},\,A_{\lambda},\,A_{\kappa},\,\mu,\,\mu^{\prime},\,m_{3}^{\prime 2},\,m_{S}^{\prime 2},\,\xi^{\prime}.$ (2.4)
In the systematic investigation of Higgs sector phenomenology, it is convenient to adopt specialized parametrizations as follows:
$H_{\rm NSM}\equiv\cos\beta{\rm Re}(H_{u}^{0})-\sin\beta{\rm Re}(H_{d}^{0}),$
$H_{\rm SM}\equiv\sin\beta{\rm Re}(H_{u}^{0})+\cos\beta{\rm Re}(H_{d}^{0}),$ (2.5)
$A_{\rm NSM}\equiv\cos\beta{\rm Im}(H_{u}^{0})-\sin\beta{\rm Im}(H_{d}^{0}).$
In the basis ($H_{\rm NSM}$, $H_{\rm SM}$, ${\rm Re[S]}$), the mass matrix of CP-even Higgs fields can be written as [28]
$\mathcal{M}^{2}_{S,11}$ $=$ $\frac{\lambda v_{s}(\sqrt{2}A_{\lambda}+\kappa v_{s}+\sqrt{2}\mu^{\prime})+2m^{2}_{3}}{\sin 2\beta}+\frac{1}{2}(2m^{2}_{Z}-\lambda^{2}v^{2})\sin^{2}2\beta,$
$\mathcal{M}^{2}_{S,12}$ $=$ $-\frac{1}{4}(2m^{2}_{Z}-\lambda^{2}v^{2})\sin 4\beta,\quad\mathcal{M}^{2}_{S,13}=-\frac{\lambda v}{\sqrt{2}}\left(A_{\lambda}+\sqrt{2}\kappa v_{s}+\mu^{\prime}\right)\cos 2\beta,$
$\mathcal{M}^{2}_{S,22}$ $=$ $m^{2}_{Z}\cos^{2}2\beta+\frac{1}{2}\lambda^{2}v^{2}\sin^{2}2\beta,$
$\mathcal{M}^{2}_{S,23}$ $=$ $\frac{\lambda v}{\sqrt{2}}\left[\left(\sqrt{2}v_{s}+2\mu\right)-(A_{\lambda}+\sqrt{2}\kappa v_{s}+\mu^{\prime})\sin 2\beta\right],$
$\mathcal{M}^{2}_{S,33}$ $=$ $\frac{\lambda v}{\sqrt{2}}\left(\frac{A_{\kappa}+\mu^{\prime}}{2\sqrt{2}}\lambda v^{2}+\frac{\kappa v_{s}}{\sqrt{2}}(A_{\kappa}+2\sqrt{2}\kappa v_{s}+3\mu^{\prime})\right)-\frac{\mu}{\sqrt{2}v_{s}}\lambda v^{2}-\frac{\sqrt{2}}{v_{s}}\xi^{\prime}.$ (2.6)
Similarly, in the basis ($A_{\rm NSM}$, ${\rm Im[S]}$), the mass matrix of CP-odd Higgs fields can be written as [28]
$\mathcal{M}^{2}_{P,11}$ $=$ $\frac{\lambda v_{s}\left(\sqrt{2}A_{\lambda}+\kappa v_{s}+\sqrt{2}\mu^{\prime}\right)+2m^{2}_{3}}{\sin 2\beta},\quad\mathcal{M}^{2}_{P,12}=\frac{\lambda v}{\sqrt{2}}\left(A_{\lambda}-\sqrt{2}\kappa v_{s}-\mu^{\prime}\right),$
$\mathcal{M}^{2}_{P,22}$ $=$ $\frac{(A_{\lambda}+2\sqrt{2}\kappa v_{s}+\mu^{\prime})\sin 2\beta}{2\sqrt{2}v_{s}}-\frac{\kappa v_{s}}{\sqrt{2}}(3A_{\kappa}+\mu^{\prime})$ (2.7)
-- $\frac{\mu}{\sqrt{2}v_{s}}\lambda v^{2}-2m^{\prime 2}_{S}-\frac{\sqrt{2}}{v_{s}}\xi^{\prime}.$
The mass eigenstates of the CP-even and CP-odd Higgs bosons arise from diagonalizing the scalar ($\mathcal{M}^{2}_{S}$) and pseudoscalar ($\mathcal{M}^{2}_{P}$) mass matrices, respectively:
$h_{i}$ $=$ $V_{h_{i}}^{\rm NSM}H_{\rm NSM}+V_{h_{i}}^{\rm SM}H_{\rm SM}+V_{h_{i}}^{\rm S}Re[S],$
$a_{j}$ $=$ $V_{P,a_{j}}^{\rm NSM}A_{\rm NSM}+V_{P,a_{j}}^{\rm S}Im[S].$ (2.8)
These correspond to three CP-even states $h_{i}$ (i=1,2,3) and two CP-odd states $a_{j}$ (j=1,2), ordered by increasing mass $m_{h_{1}}<m_{h_{2}}<m_{h_{3}}$ and $m_{a_{1}}<m_{a_{2}}$. In terms of their dominant components, the physical states are also labeled as:
  • $h$: SM-like scalar (125 GeV Higgs)
  • $H$, $A_{H}$: the CP-even and CP-odd doublet-dominated states
  • $h_{s}$, $A_{s}$: the CP-even and CP-odd singlet-dominated states
Critically, the identity of the 125 GeV Higgs depends on the mass hierarchy:
  • $h_{1}$ scenario: $h\equiv h_{1}$ with singlet-dominated $h_{s}$ heavier ($m_{h_{s}}>m_{h}$)
  • $h_{2}$ scenario: $h\equiv h_{2}$ with singlet-dominated $h_{s}$ lighter ($m_{h}>m_{h_{s}}$)
Additionally, the model predicts a pair of charged Higgs bosons:
$H^{\pm}=\cos\beta H_{u}^{\pm}+\sin\beta H_{d}^{\pm},$ (2.9)
with mass given by [8, 28]
$m^{2}_{H^{\pm}}=\frac{\lambda v_{s}\left(\sqrt{2}A_{\lambda}+\kappa v_{s}+\sqrt{2}\mu^{\prime}\right)+2m^{2}_{3}}{\sin 2\beta}+m^{2}_{W}-\frac{1}{2}\lambda^{2}v^{2}.$ (2.10)
To connect theoretical inputs with experimental observables, we replace the Lagrangian parameters $\mu,\mu^{\prime},m^{2}_{3},m^{\prime 2}_{S},$ and $\xi^{\prime}$ with physical mass parameters:
  • $m_{A}\equiv\sqrt{M^{2}_{P,11}}$: the mass of heavy MSSM-like CP-odd Higgs boson,
  • $m_{B}\equiv\sqrt{M^{2}_{S,33}}$: the mass of CP-even singlet Higgs boson,
  • $m_{C}\equiv\sqrt{M^{2}_{P,22}}$: the mass of CP-odd singlet Higgs boson,
  • $\mu_{\text{tot}}\equiv\mu_{\text{eff}}+\mu$: Higgsino mass parameter,
  • $m_{N}\equiv\frac{2\kappa}{\lambda}\mu_{\text{eff}}+\mu^{\prime}$: Singlino mass parameter.
The original Lagrangian parameters then become:
$\mu$ $=$ $\mu_{\text{tot}}-\frac{\lambda}{\sqrt{2}}v_{s},\quad\mu^{\prime}=m_{N}-\sqrt{2}\kappa v_{s},\quad m^{2}_{3}=\frac{m^{2}_{A}\sin 2\beta}{2}-\lambda v_{s}\left(\frac{\kappa v_{s}}{2}+\frac{\mu^{\prime}}{\sqrt{2}}+\frac{A_{\lambda}}{\sqrt{2}}\right),$
$\xi^{\prime}$ $=$ $\frac{v_{s}}{\sqrt{2}}\left[\frac{(A_{\lambda}+\mu^{\prime})\sin 2\beta}{2\sqrt{2}v_{s}}+\frac{\kappa v_{s}}{\sqrt{2}}(A_{\kappa}+2\sqrt{2}\kappa v_{s}+3\mu^{\prime})\right]-\frac{\mu}{\sqrt{2}v_{s}}\lambda v^{2}-m^{2}_{B},$
$m^{\prime 2}_{S}$ $=$ $\frac{1}{2}\left[m^{2}_{B}-m^{2}_{C}+\lambda\kappa\sin 2\beta v^{2}-2\sqrt{2}\kappa v_{s}(A_{\kappa}+\frac{\kappa}{\sqrt{2}}v_{s}+\mu^{\prime})\right].$ (2.11)
This reparameterization simplifies the Higgs mass matrices (Eqs. 2.6, 2.7) to:
$\mathcal{M}^{2}_{S,11}$ $=$ $m^{2}_{A}+\frac{1}{2}(2m^{2}_{Z}-\lambda^{2}v^{2})\sin^{2}2\beta,\quad\mathcal{M}^{2}_{S,12}=-\frac{1}{4}(2m^{2}_{Z}-\lambda^{2}v^{2})\sin 4\beta,$
$\mathcal{M}^{2}_{S,13}$ $=$ $-\frac{\lambda v}{\sqrt{2}}(A_{\lambda}+m_{N})\cos 2\beta,\quad\mathcal{M}^{2}_{S,22}=m^{2}_{Z}\cos^{2}2\beta+\frac{1}{2}\lambda^{2}v^{2}\sin^{2}2\beta,$
$\mathcal{M}^{2}_{S,23}$ $=$ $\frac{\lambda v}{\sqrt{2}}[2\mu_{\text{tot}}-(A_{\lambda}+m_{N})\sin 2\beta],\quad\mathcal{M}^{2}_{S,33}=m^{2}_{B},$
$\quad\mathcal{M}^{2}_{P,11}$ $=$ $m^{2}_{A},\quad\mathcal{M}^{2}_{P,22}=m^{2}_{C},\quad\mathcal{M}^{2}_{P,12}=\frac{\lambda v}{\sqrt{2}}(A_{\lambda}-m_{N}).$ (2.12)
Critically, the neutral Higgs sector is fully specified by eight parameters: $\tan\beta$, $\lambda$, $A_{\lambda}$, $m_{A}$, $m_{B}$, $m_{C}$, $m_{N}$, and $\mu_{tot}$. The remaining parameters $\kappa$, $A_{\kappa}$, and $v_{s}$ exclusively govern triple Higgs couplings [28].
In the light $h_{s}$ scenario, the scalar mass matrix element ${\cal M}^{2}_{S,23}$ is conventionally parameterized as:
${\cal M}^{2}_{S,23}=\sqrt{2}\lambda v\delta\mu_{\rm tot}$ (2.13)
where the dimensionless parameter $\delta\equiv[2\mu_{\rm tot}-(A_{\lambda}+m_{N})\sin 2\beta]/(2\mu_{\rm tot})$ quantifies the cancellation between the effective $\mu$-term and SUSY-breaking contributions. This parametrization offers significant advantages: it naturally accommodates larger $\lambda$ couplings while maintaining compatibility with LHC Higgs data through modest $\delta$ values. Consequently, we adopt $\delta$ as an input parameter instead of $A_{\lambda}$ in this work.
2.3 Trilinear Higgs couplings in the GNMSSM
For the light Higgs scenario studied in this paper, we assume a light CP-odd Higgs boson ($a_{1}$) satisfying $2m_{a_{1}}<m_{h}$, enabling the exotic decay $h\to a_{1}a_{1}$ (i.e. $h\to A_{s}A_{s}$) in both $h_{1}$ and $h_{2}$ scenarios. For the $h_{2}$ scenario, where the lightest CP-even Higgs boson $h_{1}$ is singlet-dominant, $h\to h_{1}h_{1}$ (i.e. $h\to h_{s}h_{s}$) opens when $2m_{h_{1}}<m_{h}$. Both ATLAS and CMS collaborations have performed comprehensive searches for non-SM Higgs bosons $H$, $h_{s}$, $A_{H}$ and $A_{s}$, establishing stringent exclusion bounds on key parameters including masses and relative couplings [54, 55].
The branching ratios of exotic decays $h\to a_{1}a_{1}$ and $h\to h_{1}h_{1}$ are governed by their respective trilinear Higgs couplings $C_{ha_{1}a_{1}}$($\equiv C_{hA_{s}A_{s}}$) and $C_{hh_{1}h_{1}}$ ($\equiv C_{hh_{s}h_{s}}$). In the GNMSSM, these couplings scale proportionally to vacuum expectation values ($v_{u}$, $v_{d}$, $v_{s}$) or soft SUSY-breaking trilinear terms ($A_{\lambda}$, $A_{\kappa}$). While Ref. [18] provides general expressions, we explicitly define the specific trilinear couplings relevant to our analysis below [28]:
$C_{hh_{s}h_{s}}$ $=$ $\lambda vV_{h}^{\rm SM}V_{h_{s}}^{S}V_{h_{s}}^{\rm S}(\lambda-\kappa\sin 2\beta)-\lambda\kappa vV_{h}^{\rm NSM}V_{h_{s}}^{\rm S}V_{h_{s}}^{\rm S}\cos 2\beta$ (2.14)
$+\sqrt{2}\kappa V_{h}^{\rm S}V_{h_{s}}^{\rm S}V_{h_{s}}^{\rm S}(3m_{N}+A_{\kappa})+C_{h_{s}}^{\prime}(\lambda,\kappa,\tan\beta,v_{s},A_{\lambda},m_{N})$
$C_{hA_{s}A_{s}}$ $=$ $\lambda vV_{h}^{\rm SM}(\lambda+\kappa\sin 2\beta)+\lambda\kappa vV_{h}^{\rm NSM}\cos 2\beta$ (2.15)
$+\sqrt{2}\kappa V_{h}^{\rm S}(m_{N}-A_{\kappa})+C_{A_{s}}^{\prime}(\lambda,\kappa,\tan\beta,v_{s},A_{\lambda},m_{N})$
where the last terms in the preceding equations, suppressed by Higgs mixings, depend on $\lambda,\kappa,\tan\beta,v_{s},A_{\lambda}$ and $m_{N}$.
3 Numerical Result
This section details our scanning strategy and presents numerical results. We construct the GNMSSM model file using SARAH [56, 57, 58, 59] and generate particle spectra with SPheno [60, 61, 62, 56]. For parameter space exploration, we employ the parallelized MultiNest algorithm [63], which utilizes nested sampling to efficiently identify high-likelihood regions, detect multiple solution modes, and compute Bayesian evidence.
Parameter Prior Range Parameter Prior Range
$\kappa$ Flat 0.750.75-0.75- 0.750.750.750.750.75 $\tan\beta$ Flat 555560606060
$\lambda$ Flat 000.750.750.750.75 $v_{s}/{\rm TeV}$ Flat 0.10.10.10.11.01.01.01.0
$\delta$ Flat 1.01.0-1.0- 1.01.01.01.01.0 $m_{N}/{\rm TeV}$ Flat 1.01.0-1.0- 1.01.01.01.01.0
$m_{B}/{\rm GeV}$ Flat 1111300300300300 $m_{C}/{\rm GeV}$ Flat 1.01.01.01.0300300300300
$A_{t}/{\rm TeV}$ Flat 1.01.01.01.03.03.03.03.0 $\mu_{\rm tot}/{\rm TeV}$ Flat 0.20.20.20.21.01.01.01.0
$A_{\kappa}/{\rm TeV}$ Flat 2.02.0-2.0- 2.02.02.02.02.0
Table 1: Parameter space explored in this study. All input parameters adopt flat distributions based on their unambiguous physical interpretations. Given the substantial radiative corrections induced by the third-generation squark trilinear couplings ($A_{t}$ and $A_{b}$) on the SM-like Higgs boson mass, we impose $A_{t}=A_{b}$ with their magnitudes treated as free variables. Non-critical SUSY-breaking parameters are fixed: $m_{A}=2~{}{\rm TeV}$, $M_{1}=1~{}{\rm TeV}$, $M_{2}=2~{}{\rm TeV}$, $M_{3}=3~{}{\rm TeV}$. All parameters are defined at the renormalization scale $Q_{input}=1~{}{\rm TeV}$.
3.1 Research Strategy
In the specific parameter space scan, we set the value of $n_{\text{live}}=8000$. Through multiple trial-and-error processes, we choose the GNMSSM parameter space as shown in Table 1 . The likelihood function $\mathcal{L}$ that guided our scan is influenced by relevant experimental constraints. When the parameter points satisfy the experimental limits, $\mathcal{L}=1$; otherwise, $\mathcal{L}=\exp[100]$. The basic constraints considered include:
  • The SM-like Higgs boson: In our study, we consider $h_{1}$ or $h_{2}$ as the SM-like Higgs boson $h$. The mass of $h$ has an uncertainty of 3 GeV, so it should be restricted within the range 122 GeV $\leq m_{h}\leq$ 128 GeV.
  • Higgs data fit: Assuming that $h$ corresponds to the Higgs boson discovered by the LHC, its properties should align with the measurements reported by the ATLAS and CMS collaborations at the 95% confidence level. A p-value larger than 0.05 was required, which was tested by the code HiggsSignals-2.6.2 [64, 65, 66, 67].
  • Extra Higgs searches: Additional Higgs bosons must satisfy the limits from the direct searches at LEP, Tevatron, and LHC, which are evaluated using the code HiggsBounds-5.10.2 [68, 69, 70, 71, 72].
  • $B$-physics observables: Theoretical predictions for the branching ratios of $B_{s}\to\mu^{+}\mu^{-}$ and $B\to X_{s}\gamma$ are required to be consistency with experimental measurements within $2\sigma$ [73].
Beyond the basic constraints, we incorporate constraints from HiggsTools [45], which integrates updated versions of HiggsBounds and HiggsSignals. These subpackages compare predictions against extra scalar searches (HiggsBounds) and 125 GeV Higgs measurements (HiggsSignals), respectively. Within HiggsTools, HiggsBounds thus provides direct constraints from extra scalar searches, while HiggsSignals assesses compatibility with the observed 125 GeV Higgs properties.
3.2 HiggsTools constraints
Figure 1: Surviving parameter samples after applying the basic constraints in Sec.3.1 are projected onto the plane of $\sigma(h)/\sigma_{SM}(h)Br(h\to a_{1}a_{1}\to bb\tau\tau)$ versus $m_{a_{1}}$. Blue points represent the surviving samples further constrained by HiggsTools (HT), while the black solid line denotes the ATLAS exclusion limit for the $h\to a_{1}a_{1}\to bb\tau\tau$ channel [38]. The left and right panels correspond to the $h_{1}$ and $h_{2}$ scenarios, respectively.
In the light Higgs scenario, ATLAS and CMS direct search constraints, based on decay product observations, are largely implemented in HiggsTools. To compare experimental limitations on parameter space, we project surviving samples satisfying the basic constraints in Sec. 3.1 onto two-dimensional planes with vertical axes representing $\mu_{bb\tau\tau}\equiv\sigma(h)/\sigma_{\text{SM}}(h)\times\text{Br}(h\to a_{1}a_{1}\to bb\tau\tau)$, as shown in Figure 1. Left and right panels correspond to the $h_{1}$ and $h_{2}$ scenarios, respectively. The blue points represent the surviving samples further constrained by HiggsTools. We further incorporate ATLAS constraints [38] on $h\to a_{1}a_{1}\to bb\tau\tau$ (solid line), excluding parameter space above this limit. Comparative analysis reveals that the upgraded HiggsTools imposes significantly stronger constraints than legacy HiggsSignals/HiggsBounds, excluding a large portion of parameter regions. And its exclusion power further surpasses even the latest direct search limits (solid line), which rely on single-channel analyses. After applying the constraints from HiggsTools, the values of $\mu_{bb\tau\tau}$ can reach to 1.9% in the $h_{1}$ scenario and 2.1% in the $h_{2}$ scenario, respectively.
Figure 2: Parameter samples excluded by HiggsTools in Figure 1 are color-coded based on the exclusion source. Red diamond points denote samples excluded solely by HiggsSignals (HS), orange square points represent those excluded exclusively by HiggsBounds (HB), and magenta pentagon points indicate samples excluded by both HB and HS.
To compare the constraints from HiggsBounds and HiggsSignals on the GNMSSM parameter space in the light Higgs scenario, the parameter samples excluded by HiggsTools in Figure 1 are color-coded in Figure 2 based on the exclusion source. Red diamond points denote samples excluded solely by HiggsSignals, orange square points represent those excluded exclusively by HiggsBounds, and magenta pentagon points indicate samples excluded by both tools. Figure 2 demonstrates that HiggsBounds exhibits significantly stronger exclusion power than HiggsSignals. Parameter points excluded solely by HiggsSignals (red diamonds) are nearly absent, while points excluded by HiggsSignals are almost invariably also excluded by HiggsBounds, as indicated by the dominance of magenta points.
In the $h_{1}$ scenario, parameter points with relatively large values of $\text{Br}(h\to a_{1}a_{1}\to\tau\tau bb)$ (specifically, $\text{Br}(h\to a_{1}a_{1}\to\tau\tau bb)\geq 2.5\%$) are excluded by both HiggsSignals and HiggsBounds. Points with smaller branching ratios are excluded only by HiggsBounds. This distinction arises because HiggsBounds imposes direct constraints, while HiggsSignals imposes indirect constraints on non-SM-like Higgs boson searches. Specifically speaking, an excessively large branching ratio $\text{Br}(h\to a_{1}a_{1})$, if sufficiently large, can affect the properties of the 125 GeV Higgs boson and this implies a large $\text{Br}(h\to a_{1}a_{1}\to\tau\tau bb)$, although the reverse is not necessarily true. However, small values of $\text{Br}(h\to a_{1}a_{1}\to\tau\tau bb)$ always correspond to small $\text{Br}(h\to a_{1}a_{1})$, the latter having minimal impact on the properties of the 125 GeV Higgs boson.
Figure 3: Similar to the right plot in Figure 2, here the label of the horizontal axis is $\text{Br}(h_{2}\to h_{1}h_{1}\to\tau\tau bb)$.
Consistent with the $h_{1}$ scenario, analogous characteristics are also observed in the $h_{2}$ scenario, which simultaneously manifests distinctive features. The right panel of Figure 2 shows that some parameter points with $Br(h\to a_{1}a_{1}\to\tau\tau bb)\leq 2.5\%$ (indicated by magenta points) are also excluded by HiggsSignals. This exclusion occurs because the exotic decay channel $h\to h_{1}h_{1}$ becomes kinematically accessible in this scenario, with an enhanced branching fraction. This enhancement may influence the properties of the SM-like Higgs boson and consequently affect the constraints imposed by HiggsSignals. Figure 3 plots the relationship between $Br(h\to a_{1}a_{1}\to\tau\tau bb)$ and $Br(h\to h_{1}h_{1}\to\tau\tau bb)$ specifically in the $h_{2}$ scenario. It reveals an anticorrelation between these branching ratios across the parameter space, thereby validating the proposed explanation.
Figure 4: Similar to Figure 1, this plot exhibits the relationship between the parameter $\lambda$ and $\kappa$.
Figure 5: Like Fig. 1, this plot shows the SM-like Higgs boson component in the $h_{1}$ scenario.
Figure 6: Like Fig. 1, this figure shows the $h_{2}$ component (left panel) and the $h_{1}$ component (right panel) in the $h_{2}$ scenario.
The branching ratios for the exotic decays $h\to a_{1}a_{1}$ and $h\to h_{1}h_{1}$ depend on their respective triple Higgs couplings $C_{ha_{1}a_{1}}$ and $C_{hh_{1}h_{1}}$, as defined in Eqs. (2.14) and (2.15). These couplings are primarily governed by the parameters $\lambda$, $\kappa$, and the SM-like fraction $V_{h}^{\rm SM}$ of the $h$ boson. In the $h_{2}$ scenario, $C_{hh_{1}h_{1}}$ is additionally modulated by the singlet fraction $V_{h_{s}}^{\rm S}$ of the singlet-dominated state $h_{s}$. Figure 4 displays the correlation between $\lambda$ and $\kappa$ under HiggsTools constraints, indicating $\lambda\leq 0.1$ for most viable parameter points, which are consistent with experimental preferences for minimal doublet-singlet mixing. Further, Figs. 5 and 6 reveal the Higgs boson compositions. Both $h_{1}$ and $h_{2}$ scenarios require the observed $h$ boson to be predominantly SM-like ($V_{h}^{\rm SM}\geq 0.93$) with suppressed singlet admixture ($V_{h}^{\rm S}\leq 0.32$). For the $h_{2}$ scenario, the lighter $h_{s}$ state must exhibit high singlet purity ($V_{h_{s}}^{\rm S}\geq 0.94$) and negligible SM-like components ($V_{h_{s}}^{\rm SM}\leq 0.32$).
4 Conclusion
This study systematically investigates exotic decays of the 125 GeV Higgs boson into pairs of the lightest CP-odd or CP-even Higgs bosons within the GNMSSM. Our analysis incorporates the latest constraints from ATLAS direct searches and utilizes the integrated framework of HiggsTools, which combines updated exclusion data from HiggsBounds and statistical analyses from HiggsSignals.
We first scanned the GNMSSM parameter space under fundamental constraints imposed by HiggsSignals-2.6.2 and HiggsBounds-5.10.2. The HiggsTools framework was then applied to exhaustively constrain this parameter space. A comprehensive comparison reveals that HiggsTools imposes significantly stronger constraints than previous individual versions of HiggsSignals and HiggsBounds. Furthermore, its exclusion power substantially surpasses current experimental limits derived from ATLAS direct Higgs searches, which are solely based on individual decay channel analyses.
To elucidate the impact of recent Higgs data on GNMSSM scenarios featuring a light Higgs boson, we separately analyzed the constraining effects of the subpackages HiggsSignals and HiggsBounds within HiggsTools, leading to the following key conclusions:
  • In both the $h_{1}$ and $h_{2}$ scenarios, constraints from HiggsBounds prove dominant in restricting the theoretical parameter space. Constraints from HiggsSignals and direct ATLAS searches targeting individual decay channels are comparatively weaker. Nearly all parameter points excluded by HiggsSignals or individual LHC channel analyses are also excluded by HiggsBounds. This stems from HiggsBounds providing direct constraints for non-SM Higgs searches, whereas HiggsSignals imposes indirect constraints.
  • Analyses within the $h_{2}$ scenario reveal distinct phenomenological features. Notably, HiggsSignals can exclude parameter points exhibiting suppressed exotic decay branching ratios, specifically $Br(h\to a_{1}a_{1}\to\tau\tau bb)\leq 2.5\%$. This exclusion arises due to the kinematic accessibility of the decay $h_{2}\to h_{1}h_{1}$, which features significantly enhanced branching fractions. These augmented decays modify the properties of the SM-like Higgs boson, thereby increasing the sensitivity of HiggsSignals’ indirect constraints in otherwise kinematically suppressed regions.
  • Under the combined constraints of HiggsTools, both the $h_{1}$ and $h_{2}$ scenarios require the observed $h$ boson to be predominantly SM-like ($V_{h}^{\rm SM}\geq 0.93$) with minimal singlet admixture ($V_{h}^{\rm S}\leq 0.32$). In the $h_{2}$ scenario, the lighter $h_{s}$ state must exhibit high singlet purity ($V_{h_{s}}^{\rm S}\geq 0.94$) and negligible SM-like components ($V_{h_{s}}^{\rm SM}\leq 0.32$).
Crucially, despite the rigorous constraints imposed by current Higgs data on light Higgs scenarios, phenomenologically viable regions of parameter space persist. These regions offer promising prospects for future exploration of light Higgs phenomenology.
Acknowledgement
We sincerely thank Prof.Junjie Cao for helpful discussions.