### v3

# Constraining the light Higgs bosons in the GNMSSM with recent Higgs data

Abstract
The search for light scalar and pseudoscalar particles not only provides a promising avenue for probing physics beyond the Standard Model, but also enables a strong first-order electroweak phase transition, thereby offering a viable mechanism to explain the observed baryon asymmetry of the universe. In this study, we investigate exotic decay channels of the 125 GeV Standard-Model-like Higgs boson into pairs of light CP-odd ($a_{s}$) or CP-even ($h_{s}$) Higgs bosons within the framework of the Generalized Next-to-Minimal Supersymmetric Standard Model (GNMSSM). A comprehensive parameter space scan is performed using the MultiNest algorithm, incorporating constraints from HiggsSignals-2.6.2, HiggsBounds-5.10.2, and ATLAS experimental searches, under two distinct scenarios(one in which the observed Higgs boson corresponds to the lighter CP-even state ($h_{1}=h_{\text{SM}}$), and another where it is identified with the heavier state ($h_{2}=h_{\text{SM}}$) ). Our results demonstrate that HiggsBounds imposes the most stringent exclusion limits due to its sensitivity to direct searches for non-Standard Model Higgs bosons. In contrast, HiggsSignals can also exclude regions with suppressed exotic branching ratios—such as $Br(h\to a_{s}a_{s}\to\tau\tau bb)\leq 2.5\%$ in the $h_{2}$-SM scenario—by virtue of its ability to probe indirect deviations arising from kinematically enhanced $h_{2}\to h_{s}h_{s}$ decays. Under combined constraints from HiggsTools, the observed Higgs boson must retain at least 93% SM-like component ($V_{h}^{\text{SM}}\geq 0.93$) with no more than 32% singlet admixture ($V_{h}^{\text{S}}\leq 0.32$); in the $h_{2}$ case, the lighter scalar $h_{s}$ exhibits high singlet purity ($V_{h_{s}}^{\text{S}}\geq 0.94$). Furthermore, dark matter (DM) phenomenology indicates that Singlino- or Higgsino-dominated DM is viable in the $h_{1}$ scenario, with dominant annihilation channels including $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to h_{s}A_{s}$ for Singlino-like DM and chargino co-annihilation for Higgsino-like DM, whereas the $h_{2}$ scenario favors Higgsino-dominated DM.
1 Introduction
The discovery of the 125-GeV Higgs boson at the Large Hadron Collider (LHC) [1, 2] marked a milestone in particle physics because it confirmed the mechanism of electroweak symmetry breaking (EWSB) that the Standard Model (SM) describes. However, the SM leaves several fundamental problems unsolved, such as the neutrino masses, hierarchy problem, and dark matter (DM) problem. These unresolved issues have motivated exploration of physics beyond the Standard Model (BSM), where extended Higgs sectors and new symmetries often predict additional scalar states.
Supersymmetry (SUSY) is a BSM candidate that offers solutions to address the hierarchy problem and enable gauge coupling unification. The Minimum Supersymmetric Standard Model (MSSM) [3, 4, 5, 6], which is the most economical SUSY realization, has been extensively studied. However, it is challenged by DM constraints, the $\mu$ problem, the little hierarchy problem, and other issues. As a natural and minimal extension, the Next-to-Minimal Supersymmetric Standard Model (NMSSM) [7, 8, 9] introduces a gauge-singlet Higgs superfield. Its scalar component acquires a non-zero vacuum expectation value (vev), which generates an effective $\mu$-term. Mixing between singlet and doublet superfields produces a new CP-even Higgs ($h_{s}$) that mixes with MSSM-like states. This mixing can suppress the lightest Higgs mass while enabling the next-to-lightest Higgs to match the observed 125-GeV signal [10, 11, 12, 13, 14, 15, 16]. This scenario alleviates fine-tuning and offers rich phenomenology, including Higgs-to-Higgs decays [17, 18, 19, 20, 21, 22, 23, 24] or exotic decays that involve neutralinos [25].
Over the past decade, the NMSSM with the $Z_{3}$-symmetric ($Z_{3}$-invariant NMSSM) [7, 8, 9] has drawn significant interest as a simple extension of the MSSM. However, LHC searches for supersymmetric particles and tighter limits from dark matter experiments such as XENON-nT [26], PandaX-4T [27, 28], and LUX-ZEPLIN (LZ) [29, 30] have severely constrained its parameter space. Searches for light Higgs bosons, electroweakinos, and extra scalars have ruled out much of the natural parameter region, especially where the 125 GeV Higgs mass, B-physics, and dark matter density are simultaneously satisfied [31, 32]. For large $\lambda$, predicted spin-independent scattering cross sections approach current detection limits [31, 32]; only small $\lambda$ or phase cancellations allow evasion, but small $\lambda$ complicates achieving the correct Higgs mass without fine-tuning [31, 32]. Thus, the $Z_{3}$-invariant NMSSM survives only in a narrow and fine-tuned region, challenging its original goal of solving the naturalness problem.
To address the aforementioned theoretical limitations of the $Z_{3}$-invariant NMSSM regarding the Higgs boson mass, the superparticle mass spectrum, and its connection to dark matter, researchers have proposed various extended frameworks, among which the General NMSSM (GNMSSM) [33, 34, 35, 36, 37, 38] stands out as a representative example. By adding extra singlet couplings or $Z_{3}$-symmetry-breaking terms, the GNMSSM enhances the flexibility in the Higgs mass matrix and enables lighter CP-even and CP-odd states, which are consistent with LHC data. Moreover, within the framework of the GNMSSM, the lightest supersymmetric particle (LSP) is a dark matter (DM) candidate. Its annihilation mainly proceeds through s-channel exchange of light CP-even and CP-odd Higgs bosons. The coupling strength determines both the dark matter relic abundance from the early universe and the elastic scattering cross-section with nucleons in direct detection experiments. This link connects collider physics and dark matter phenomenology, allowing constraints from high-energy and astrophysical observations to complement each other.
Couplings between these light scalar particles and standard model fermions or gauge bosons are dynamically suppressed, leading to very low production rates in diphoton, ZZ, or WW channels. As a result, these particles are difficult to detect with conventional searches. However, they can still appear indirectly through hidden-sector effects, especially in exotic Higgs decays. In certain regions, the 125 GeV Higgs boson h decays into pairs of light scalars (e.g., $h\to h_{s}h_{s}$ or $a_{s}a_{s}$ ), each of which then decays into b-jets, $\tau$-lepton pairs, or other standard model states. Although these cascade decays have small branching ratios, their low backgrounds make them sensitive probes for new physics. Therefore, ATLAS and CMS have systematically searched for such signatures across multiple LHC runs, targeting final states such as $4b$ [39], $\mu\mu bb$ [40, 41, 42], $\tau\tau bb$ [43, 44], $\gamma\gamma\tau\tau$ [45], and $4\mu$ [46, 47, 48, 49, 50].
This study focuses on the $\tau\tau bb$ final state [44]. Compared to the $4b$ [39], this signature offers enhanced background suppression by exploiting the distinctive decay topology of the $\tau$ lepton, while retaining high signal efficiency. We use the HiggsTools package[51] to constrain the model parameter space, identifying regions consistent with direct search limits and Higgs property measurements. Based on these constraints, we study $h\to h_{s}h_{s}/a_{s}a_{s}$ exotic decays in the GNMSSM framework and evaluate their observability. We have also conducted a preliminary analysis of the properties of dark matter within the allowed parameter region.
This paper is organized as follows. Section 2 introduces the GNMSSM framework and describes in detail its Higgs sector and the key trilinear Higgs couplings. Section 3 outlines our scanning strategy and the corresponding numerical results. Section 4 summarizes our conclusions.
2 Theoretical preliminaries
2.1 Basics of the GNMSSM
The GNMSSM extends the MSSM by introducing a gauge-singlet superfield $\hat{S}$, which carries neither baryon nor lepton number. Thus, the Higgs sector comprises the standard $SU(2)_{L}$ doublets $\hat{H}_{u}=(\hat{H}^{+}_{u},\hat{H}^{0}_{u})$ and $\hat{H}_{d}=(\hat{H}^{0}_{d},\hat{H}^{-}_{d})$ and the singlet $\hat{S}$. Its gauge-invariant superpotential is as follows  [8]:
$W_{\rm GNMSSM}=W_{\text{Yukawa}}+\lambda\hat{S}\hat{H_{u}}\cdot\hat{H_{d}}+\frac{\kappa}{3}\hat{S}^{3}+\mu\hat{H_{u}}\cdot\hat{H_{d}}+\frac{1}{2}\mu^{\prime}\hat{S}^{2}+\xi\hat{S}$, (2.1)
where $W_{\rm Yukawa}$ contains the MSSM quark and lepton Yukawa interactions. The dimensionless couplings $\lambda$ and $\kappa$ govern the Higgs interactions, which are analogous to the $Z_{3}$-NMSSM. The GNMSSM incorporates explicit $Z_{3}$-symmetry-breaking via parameters $\mu$, $\mu^{\prime}$ and $\xi$. These terms resolve the inherent tadpole problem [52, 8] and cosmological domain-wall problem [53, 54, 55] in the $Z_{3}$-invariant NMSSM. Parameter $\xi$ can be eliminated by singlet field redefinition  [56] and enables us to set $\xi=0$ without loss of generality. $\mu$ and $\mu^{\prime}$ arise from the spontaneous breaking of discrete $R$-symmetries ($\mathbb{Z}^{R}_{4}$ or $\mathbb{Z}^{R}_{8}$) at high scales [53, 57, 58, 56, 59]. These explicit $Z_{3}$-breaking terms modify the neutral Higgs mass spectrum and generate markedly richer phenomenology than the $Z_{3}$-NMSSM and MSSM.
2.2 Higgs sector of the GNMSSM
The soft SUSY-breaking terms in the Higgs sector of the GNMSSM can be expressed as:
$-\mathcal{L}_{soft}=$ $\Bigg[\lambda A_{\lambda}SH_{u}\cdot H_{d}+\frac{1}{3}\kappa A_{\kappa}S^{3}+m_{3}^{2}H_{u}\cdot H_{d}+\frac{1}{2}{m_{S}^{\prime}}^{2}S^{2}+\xi^{\prime}S+h.c.\Bigg]$ (2.2)
$+m^{2}_{H_{u}}|H_{u}|^{2}+m^{2}_{H_{d}}|H_{d}|^{2}+m^{2}_{S}|S|^{2},$
where $H_{u},H_{d},$ and $S$ are scalar components of the Higgs superfields, and $m^{2}_{H_{u}},m^{2}_{H_{d}},$ and $m^{2}_{S}$ are their supersymmetry-breaking masses. After the electroweak symmetry breaking, the neutral components of the Higgs fields develop non-vanishing vevs:
$\langle H^{0}_{u}\rangle=\frac{v_{u}}{\sqrt{2}},\quad\langle H^{0}_{d}\rangle=\frac{v_{d}}{\sqrt{2}},\quad\langle S\rangle=\frac{v_{s}}{\sqrt{2}},$ (2.3)
with $v=\sqrt{v_{u}^{2}+v_{d}^{2}}\simeq 246\,\text{GeV}$. Then, the Higgs sector is characterized by 11 independent physical parameters:
$\tan\beta\equiv\frac{v_{u}}{v_{d}},\,\lambda,\,\kappa,\,v_{s},\,A_{\lambda},\,A_{\kappa},\,\mu,\,\mu^{\prime},\,m_{3}^{\prime 2},\,m_{S}^{\prime 2},\,\xi^{\prime}.$ (2.4)
In the investigation of the Higgs sector phenomenology, it is useful to adopt specialized parametrizations as follows:
$H_{\rm NSM}\equiv\cos\beta{\rm Re}(H_{u}^{0})-\sin\beta{\rm Re}(H_{d}^{0}),$
$H_{\rm SM}\equiv\sin\beta{\rm Re}(H_{u}^{0})+\cos\beta{\rm Re}(H_{d}^{0}),$ (2.5)
$A_{\rm NSM}\equiv\cos\beta{\rm Im}(H_{u}^{0})-\sin\beta{\rm Im}(H_{d}^{0}).$
In the basis of ($H_{\rm NSM}$, $H_{\rm SM}$, $\rm Re[S]$}), the mass matrix of CP-even Higgs fields can be written as: [35]
$\mathcal{M}^{2}_{S,11}$ $=$ $\frac{\lambda v_{s}(\sqrt{2}A_{\lambda}+\kappa v_{s}+\sqrt{2}\mu^{\prime})+2m^{2}_{3}}{\sin 2\beta}+\frac{1}{2}(2m^{2}_{Z}-\lambda^{2}v^{2})\sin^{2}2\beta,$
$\mathcal{M}^{2}_{S,12}$ $=$ $-\frac{1}{4}(2m^{2}_{Z}-\lambda^{2}v^{2})\sin 4\beta,\quad\mathcal{M}^{2}_{S,13}=-\frac{\lambda v}{\sqrt{2}}\left(A_{\lambda}+\sqrt{2}\kappa v_{s}+\mu^{\prime}\right)\cos 2\beta,$
$\mathcal{M}^{2}_{S,22}$ $=$ $m^{2}_{Z}\cos^{2}2\beta+\frac{1}{2}\lambda^{2}v^{2}\sin^{2}2\beta,$
$\mathcal{M}^{2}_{S,23}$ $=$ $\frac{\lambda v}{\sqrt{2}}\left[\left(\sqrt{2}v_{s}+2\mu\right)-(A_{\lambda}+\sqrt{2}\kappa v_{s}+\mu^{\prime})\sin 2\beta\right],$
$\mathcal{M}^{2}_{S,33}$ $=$ $\frac{\lambda v}{\sqrt{2}}\left(\frac{A_{\kappa}+\mu^{\prime}}{2\sqrt{2}}\lambda v^{2}+\frac{\kappa v_{s}}{\sqrt{2}}(A_{\kappa}+2\sqrt{2}\kappa v_{s}+3\mu^{\prime})\right)-\frac{\mu}{\sqrt{2}v_{s}}\lambda v^{2}-\frac{\sqrt{2}}{v_{s}}\xi^{\prime}.$ (2.6)
Similarly, based on ($A_{\rm NSM}$, $\rm Im[S]$}), the mass matrix of CP-odd Higgs fields can be written as [35]
$\mathcal{M}^{2}_{P,11}$ $=$ $\frac{\lambda v_{s}\left(\sqrt{2}A_{\lambda}+\kappa v_{s}+\sqrt{2}\mu^{\prime}\right)+2m^{2}_{3}}{\sin 2\beta},\quad\mathcal{M}^{2}_{P,12}=\frac{\lambda v}{\sqrt{2}}\left(A_{\lambda}-\sqrt{2}\kappa v_{s}-\mu^{\prime}\right),$
$\mathcal{M}^{2}_{P,22}$ $=$ $\frac{(A_{\lambda}+2\sqrt{2}\kappa v_{s}+\mu^{\prime})\sin 2\beta}{2\sqrt{2}v_{s}}-\frac{\kappa v_{s}}{\sqrt{2}}(3A_{\kappa}+\mu^{\prime})$ (2.7)
- $\frac{\mu}{\sqrt{2}v_{s}}\lambda v^{2}-2m^{\prime 2}_{S}-\frac{\sqrt{2}}{v_{s}}\xi^{\prime}.$
The mass eigenstates of the CP-even and CP-odd Higgs bosons arise from diagonalizing the scalar ($\mathcal{M}^{2}_{S}$) and pseudoscalar ($\mathcal{M}^{2}_{P}$) mass matrices, respectively:
$h_{i}$ $=$ $V_{h_{i}}^{\rm NSM}H_{\rm NSM}+V_{h_{i}}^{\rm SM}H_{\rm SM}+V_{h_{i}}^{\rm S}Re[S],$
$a_{j}$ $=$ $V_{P,a_{j}}^{\rm NSM}A_{\rm NSM}+V_{P,a_{j}}^{\rm S}Im[S].$ (2.8)
These mass matrices correspond to three CP-even states $h_{i}$ (i=1,2,3) and two CP-odd states $a_{j}$ (j=1,2), which are ordered by increasing mass: $m_{h_{1}}<m_{h_{2}}<m_{h_{3}}$ and $m_{a_{1}}<m_{a_{2}}$. In terms of dominant components, the physical states are labeled as follows:
  • $h$: SM-like scalar (125-GeV Higgs)
  • $H$, $A_{H}$: CP-even and CP-odd doublet-dominated states
  • $h_{s}$, $A_{s}$: CP-even and CP-odd singlet-dominated states
The identity of the 125-GeV Higgs depends on the mass hierarchy:
  • $h_{1}$ scenario: $h\equiv h_{1}$ with singlet-dominated $h_{s}$ heavier ($m_{h_{s}}>m_{h}$)
  • $h_{2}$ scenario: $h\equiv h_{2}$ with singlet-dominated $h_{s}$ lighter ($m_{h}>m_{h_{s}}$)
In addition, the model predicts a pair of charged Higgs bosons:
$H^{\pm}=\cos\beta H_{u}^{\pm}+\sin\beta H_{d}^{\pm},$ (2.9)
with the following mass: [8, 35]
$m^{2}_{H^{\pm}}=\frac{\lambda v_{s}\left(\sqrt{2}A_{\lambda}+\kappa v_{s}+\sqrt{2}\mu^{\prime}\right)+2m^{2}_{3}}{\sin 2\beta}+m^{2}_{W}-\frac{1}{2}\lambda^{2}v^{2}.$ (2.10)
To connect the theoretical inputs with experimental observables, we replaced the Lagrangian parameters $\mu,\mu^{\prime},m^{2}_{3},m^{\prime 2}_{S},$ and $\xi^{\prime}$ with physical mass parameters:
  • $m_{A}\equiv\sqrt{M^{2}_{P,11}}$: mass of a heavy MSSM-like CP-odd Higgs boson,
  • $m_{B}\equiv\sqrt{M^{2}_{S,33}}$: mass of a CP-even singlet Higgs boson,
  • $m_{C}\equiv\sqrt{M^{2}_{P,22}}$: mass of a CP-odd singlet Higgs boson,
  • $\mu_{\text{tot}}\equiv\mu_{\text{eff}}+\mu$: Higgsino mass parameter,
  • $m_{N}\equiv\frac{2\kappa}{\lambda}\mu_{\text{eff}}+\mu^{\prime}$: Singlino mass parameter.
Then, the original Lagrangian parameters become:
$\mu$ $=$ $\mu_{\text{tot}}-\frac{\lambda}{\sqrt{2}}v_{s},\quad\mu^{\prime}=m_{N}-\sqrt{2}\kappa v_{s},\quad m^{2}_{3}=\frac{m^{2}_{A}\sin 2\beta}{2}-\lambda v_{s}\left(\frac{\kappa v_{s}}{2}+\frac{\mu^{\prime}}{\sqrt{2}}+\frac{A_{\lambda}}{\sqrt{2}}\right),$
$\xi^{\prime}$ $=$ $\frac{v_{s}}{\sqrt{2}}\left[\frac{(A_{\lambda}+\mu^{\prime})\sin 2\beta}{2\sqrt{2}v_{s}}+\frac{\kappa v_{s}}{\sqrt{2}}(A_{\kappa}+2\sqrt{2}\kappa v_{s}+3\mu^{\prime})\right]-\frac{\mu}{\sqrt{2}v_{s}}\lambda v^{2}-m^{2}_{B},$
$m^{\prime 2}_{S}$ $=$ $\frac{1}{2}\left[m^{2}_{B}-m^{2}_{C}+\lambda\kappa\sin 2\beta v^{2}-2\sqrt{2}\kappa v_{s}(A_{\kappa}+\frac{\kappa}{\sqrt{2}}v_{s}+\mu^{\prime})\right].$ (2.11)
This reparameterization simplifies the Higgs mass matrices (Eqs. 2.6, 2.7) to:
$\mathcal{M}^{2}_{S,11}$ $=$ $m^{2}_{A}+\frac{1}{2}(2m^{2}_{Z}-\lambda^{2}v^{2})\sin^{2}2\beta,\quad\mathcal{M}^{2}_{S,12}=-\frac{1}{4}(2m^{2}_{Z}-\lambda^{2}v^{2})\sin 4\beta,$
$\mathcal{M}^{2}_{S,13}$ $=$ $-\frac{\lambda v}{\sqrt{2}}(A_{\lambda}+m_{N})\cos 2\beta,\quad\mathcal{M}^{2}_{S,22}=m^{2}_{Z}\cos^{2}2\beta+\frac{1}{2}\lambda^{2}v^{2}\sin^{2}2\beta,$
$\mathcal{M}^{2}_{S,23}$ $=$ $\frac{\lambda v}{\sqrt{2}}[2\mu_{\text{tot}}-(A_{\lambda}+m_{N})\sin 2\beta],\quad\mathcal{M}^{2}_{S,33}=m^{2}_{B},$
$\quad\mathcal{M}^{2}_{P,11}$ $=$ $m^{2}_{A},\quad\mathcal{M}^{2}_{P,22}=m^{2}_{C},\quad\mathcal{M}^{2}_{P,12}=\frac{\lambda v}{\sqrt{2}}(A_{\lambda}-m_{N}).$ (2.12)
The neutral Higgs sector is fully specified by eight parameters: $\tan\beta$, $\lambda$, $A_{\lambda}$, $m_{A}$, $m_{B}$, $m_{C}$, $m_{N}$, and $\mu_{tot}$. The remaining parameters ($\kappa$, $A_{\kappa}$, and $v_{s}$) exclusively govern triple Higgs couplings [35].
In the light $h_{s}$ scenario, the scalar mass matrix element ${\cal M}^{2}_{S,23}$ is parameterized as:
${\cal M}^{2}_{S,23}=\sqrt{2}\lambda v\delta\mu_{\rm tot}$ (2.13)
where the dimensionless parameter $\delta\equiv[2\mu_{\rm tot}-(A_{\lambda}+m_{N})\sin 2\beta]/(2\mu_{\rm tot})$ quantifies the cancellation between effective $\mu$-term and SUSY-breaking contributions. This parametrization offers advantages: it naturally accommodates larger $\lambda$ couplings while remaining compatible with LHC Higgs data through modest $\delta$ values. Thus, we used $\delta$ as an input parameter instead of $A_{\lambda}$ in this work.
In the scenario with very heavy charged Higgs bosons, approximate expressions can be derived for the singlet-dominated states and their mixing parameters [60]:
$m_{h_{s}}^{2}$ $\simeq$ $m_{B}^{2}-\frac{{\cal M}^{4}_{S,13}}{m_{A}^{2}-m_{B}^{2}},\quad m_{A_{s}}^{2}\simeq m_{C}^{2}-\frac{{\cal M}^{4}_{P,12}}{m_{A}^{2}-m_{C}^{2}},\quad\frac{V_{P,A_{s}}^{\rm NSM}}{V_{P,A_{s}}^{\rm S}}=\frac{{\cal M}^{2}_{P,12}}{m_{A_{s}}^{2}-m_{A}^{2}}\simeq 0,$
$\frac{V_{h}^{\rm S}}{V_{h}^{\rm SM}}$ $\simeq$ $\frac{{\cal M}^{2}_{S,23}}{m_{h}^{2}-m_{B}^{2}},\quad V_{h}^{\rm NSM}\sim 0,\quad V_{h}^{\rm SM}\simeq\sqrt{1-\left(\frac{V_{h}^{\rm S}}{V_{h}^{\rm SM}}\right)^{2}}\sim 1,$
$\frac{V_{h_{s}}^{\rm SM}}{V_{h_{s}}^{\rm S}}$ $\simeq$ $\frac{{\cal M}^{2}_{S,23}}{m_{h_{s}}^{2}-m_{h}^{2}},\quad V_{h_{s}}^{\rm NSM}\sim 0,\quad V_{h_{s}}^{\rm S}\simeq\sqrt{1-\left(\frac{V_{h_{s}}^{\rm SM}}{V_{h_{s}}^{\rm S}}\right)^{2}}\sim 1.$ (2.14)
These expressions indicate that $V_{h}^{S}\simeq-V_{h_{s}}^{\rm SM}$ and that all scale with $\lambda$. In the limit of $\lambda\to 0$, the singlet fields decouple from the doublet Higgs fields, and the masses $m_{B}$, $m_{C}$, and $m_{N}$ can be treated as the physical particle masses with high accuracy. Moreover, the singlet masses are independent and may assume small values, as they are only weakly constrained by current experimental data.
2.3 Neutralino sector of the GNMSSM
[…]
$M_{\tilde{\chi}^{0}}=\left(\begin{array}[]{ccccc}M_{1}&0&-m_{Z}\sin\theta_{W}\cos\beta&m_{Z}\sin\theta_{W}\sin\beta&0\\ &M_{2}&m_{Z}\cos\theta_{W}\cos\beta&-m_{Z}\cos\theta_{W}\sin\beta&0\\ &&0&-\mu_{\text{tot}}&-\frac{1}{\sqrt{2}}\lambda v\sin\beta\\ &&&0&-\frac{1}{\sqrt{2}}\lambda v\cos\beta\\ &&&&m_{\text{N}}\end{array}\right),$ (2.15)
where $M_{1}$ and $M_{2}$ denote the gaugino soft-breaking mass parameters, while $s_{W}\equiv\sin\theta_{W}$ and $c_{W}\equiv\cos\theta_{W}$. By diagonalization the mass matrix using the rotation matrix $N$, the physical neutralino mass eigenstates $\tilde{\chi}^{0}_{i}$ are obtained as follows:
$\tilde{\chi}_{i}^{0}=N_{i1}\psi^{0}_{1}+N_{i2}\psi^{0}_{2}+N_{i3}\psi^{0}_{3}+N_{i4}\psi^{0}_{4}+N_{i5}\psi^{0}_{5},$ (2.16)
where the index i=1,2,,5i=1,2,...,5, arranged in ascending order of mass. The lightest neutralino, $\tilde{\chi}_{1}^{0}$, can serve as the DM candidate. In this state, $N_{13}^{2}+N_{14}^{2}$ and $N_{15}^{2}$ correspond to the Higgsino and Singlino components of the physical state $\tilde{\chi}_{1}^{0}$, respectively. The lightest neutralino $\tilde{\chi}_{1}^{0}$ is classified as Higgsino-dominated DM if $(N_{13}^{2}+N_{14}^{2})>0.5$ and Singlino-dominated DM if $N_{15}^{2}>0.5$. In scenarios where gauginos are significantly heavier and the condition $\mu_{tot}^{2}-m_{N}^{2}>>\lambda^{2}v^{2}$ holds, the mass of the Singlino-dominated DM can be approximately given by [35]
$m_{\tilde{\chi}_{1}^{0}}\simeq m_{N}+\frac{\lambda^{2}v^{2}(m_{\tilde{\chi}_{1}^{0}}-\mu_{tot}\sin 2\beta)}{2(m_{\tilde{\chi}_{1}^{0}}^{2}-\mu_{tot}^{2})}\simeq m_{N}$ (2.17)
2.4 Trilinear Higgs couplings in the GNMSSM
For the light Higgs scenario in this study, we assumed a light CP-odd Higgs boson ($a_{1}$) that satisfied $2m_{a_{1}}<m_{h}$ to enable the exotic decay $h\to a_{1}a_{1}$ (i.e. $h\to A_{s}A_{s}$) in both $h_{1}$ and $h_{2}$ scenarios. For the $h_{2}$ scenario, where the lightest CP-even Higgs boson $h_{1}$ is singlet-dominant, $h\to h_{1}h_{1}$ (i.e., $h\to h_{s}h_{s}$) opens when $2m_{h_{1}}<m_{h}$. Both ATLAS and CMS collaborations have searched for non-SM Higgs bosons $H$, $h_{s}$, $A_{H}$ and $A_{s}$, which established stringent exclusion bounds on key parameters, including the masses and relative couplings [61, 62].
The branching ratios of exotic decays $h\to a_{1}a_{1}$ and $h\to h_{1}h_{1}$ are governed by their respective trilinear Higgs couplings $C_{ha_{1}a_{1}}$($\equiv C_{hA_{s}A_{s}}$) and $C_{hh_{1}h_{1}}$ ($\equiv C_{hh_{s}h_{s}}$). In the GNMSSM, these couplings proportionally scale to vacuum expectation values ($v_{u}$, $v_{d}$, $v_{s}$) or soft SUSY-breaking trilinear terms ($A_{\lambda}$, $A_{\kappa}$). Ref. [18] provides general expressions, but we explicitly define specific trilinear couplings that are relevant to our analysis below [35]:
$C_{hh_{s}h_{s}}$ $=$ $\lambda vV_{h}^{\rm SM}V_{h_{s}}^{S}V_{h_{s}}^{\rm S}(\lambda-\kappa\sin 2\beta)-\lambda\kappa vV_{h}^{\rm NSM}V_{h_{s}}^{\rm S}V_{h_{s}}^{\rm S}\cos 2\beta$ (2.18)
$+\sqrt{2}\kappa V_{h}^{\rm S}V_{h_{s}}^{\rm S}V_{h_{s}}^{\rm S}(3m_{N}+A_{\kappa})+C_{h_{s}}^{\prime}(\lambda,\kappa,\tan\beta,v_{s},A_{\lambda},m_{N})$
$C_{hA_{s}A_{s}}$ $=$ $\lambda vV_{h}^{\rm SM}(\lambda+\kappa\sin 2\beta)+\lambda\kappa vV_{h}^{\rm NSM}\cos 2\beta$ (2.19)
$+\sqrt{2}\kappa V_{h}^{\rm S}(m_{N}-A_{\kappa})+C_{A_{s}}^{\prime}(\lambda,\kappa,\tan\beta,v_{s},A_{\lambda},m_{N})$
where the last terms in the preceding equations, which are suppressed by Higgs mixings, depend on $\lambda,\kappa,\tan\beta,v_{s},A_{\lambda}$ and $m_{N}$.
3 Numerical result
We constructed the GNMSSM model file using SARAH [63, 64, 65, 66] and generate particle spectra with SPheno [67, 68, 69, 63]. To explore the parameter space, we used the parallelized MultiNest algorithm [70], which uses nested sampling to identify high-likelihood regions, detect multiple solution modes, and compute Bayesian evidence.
Parameter Prior Range Parameter Prior Range
$\kappa$ Flat 0.75-0.750.750.75 $\tan\beta$ Flat 556060
$\lambda$ Flat 00.750.75 $v_{s}/{\rm TeV}$ Flat 0.10.11.01.0
$\delta$ Flat 1.0-1.01.01.0 $m_{N}/{\rm TeV}$ Flat 1.0-1.01.01.0
$m_{B}/{\rm GeV}$ Flat 11300300 $m_{C}/{\rm GeV}$ Flat 1.01.0300300
$A_{t}/{\rm TeV}$ Flat 1.01.03.03.0 $\mu_{\rm tot}/{\rm TeV}$ Flat 0.20.21.01.0
$A_{\kappa}/{\rm TeV}$ Flat 2.0-2.02.02.0
Table 1: Parameter space explored in this study. All input parameters used flat distributions based on their unambiguous physical interpretations. Considering the substantial radiative corrections that the third-generation squark trilinear couplings ($A_{t}$ and $A_{b}$) impose on the SM-like Higgs boson mass, we imposed $A_{t}=A_{b}$, where their magnitudes were treated as free variables. Non-critical SUSY-breaking parameters were fixed: $m_{A}=2~{\rm TeV}$, $M_{1}=1~{\rm TeV}$, $M_{2}=2~{\rm TeV}$, $M_{3}=3~{\rm TeV}$. All parameters were defined at the renormalization scale $Q_{input}=1~{\rm TeV}$.
3.1 Research strategy
In the specific parameter space scan, we set the value of $n_{\text{live}}=8000$. Through trial-and-error processes, we selected the GNMSSM parameter space as shown in Table 1. The likelihood function $\mathcal{L}$ that guided our scan is influenced by experimental constraints. When the parameter points satisfy the experimental limits, $\mathcal{L}=1$; otherwise, $\mathcal{L}=\exp[100]$. The basic constraints that we considered include:
  • The SM-like Higgs boson: In our study, we considered $h_{1}$ or $h_{2}$ as the SM-like Higgs boson $h$. The mass of $h$ has an uncertainty of 3 GeV, so it should be restricted in the range of 122 GeV $\leq m_{h}\leq$ 128 GeV.
  • Higgs data fit: Assuming that $h$ corresponds to the Higgs boson discovered by the LHC, its properties should align with the measurements reported by the ATLAS and CMS collaborations at the 95% confidence level. A p-value higher than 0.05 was required, which was tested by the code HiggsSignals-2.6.2 [71, 72, 73, 74].
  • Extra Higgs searches: Additional Higgs bosons must satisfy the limits from the direct searches at the LEP, Tevatron, and LHC, which were evaluated using the code HiggsBounds-5.10.2 [75, 76, 77, 78, 79].
  • $B$-physics observables: Theoretical predictions for the branching ratios of $B_{s}\to\mu^{+}\mu^{-}$ and $B\to X_{s}\gamma$ are required to be consistent with experimental measurements within $2\sigma$ [80].
Beyond the basic constraints, we incorporate constraints from HiggsTools [51], which integrates the updated versions of HiggsBounds and HiggsSignals. These subpackages compare predictions against extra scalar searches (HiggsBounds) and 125-GeV Higgs measurements (HiggsSignals), respectively. In HiggsTools, HiggsBounds provides direct constraints from extra scalar searches, whereas HiggsSignals assesses the compatibility with the observed 125-GeV Higgs properties.
3.2 HiggsTools constraints
Figure 1: Surviving parameter samples after we applied the basic constraints in Sec. 3.1 were projected onto the plane of $\sigma(h)/\sigma_{SM}(h)Br(h\to a_{1}a_{1}\to bb\tau\tau)$ versus $m_{a_{1}}$. The blue points represent the surviving samples after further constraints by HiggsTools (HT); the black solid line denotes the ATLAS exclusion limit for the $h\to a_{1}a_{1}\to bb\tau\tau$ channel [44]. The left and right panels correspond to the $h_{1}$ and $h_{2}$ scenarios, respectively.
In the light Higgs scenario, the direct search constraints of ATLAS and CMS, which are based on decay product observations, were largely implemented in HiggsTools. To compare experimental limitations on the parameter space, we projected the surviving samples that satisfied the basic constraints in Sec. 3.1 onto two-dimensional planes, where the vertical axes represent $\mu_{bb\tau\tau}\equiv\sigma(h)/\sigma_{\text{SM}}(h)\times\text{Br}(h\to a_{1}a_{1}\to bb\tau\tau)$, as shown in Figure 1. The left and right panels correspond to the $h_{1}$ and $h_{2}$ scenarios, respectively. The blue points represent the surviving samples after further constraints of HiggsTools. We incorporated the ATLAS constraints [44] on $h\to a_{1}a_{1}\to bb\tau\tau$ (solid line) and excluded the parameter space above this limit. Comparative analysis reveals that the upgraded HiggsTools imposes stronger constraints than the legacy HiggsSignals/HiggsBounds. Approximately 96.7% (95.1%) of the parameter points were excluded in the $h_{1}$ ($h_{2}$) scenario. In addition, its exclusion power surpasses even the latest direct search limits (solid line), which rely on single-channel analyses. After the constraints from HiggsTools had been applied, the values of $\mu_{bb\tau\tau}$ could reach 1.9% in the $h_{1}$ scenario and 2.1% in the $h_{2}$ scenario.
Figure 2: Parameter samples excluded by HiggsTools in Figure 1 are color-coded based on the exclusion source. The red diamond points denote samples that only HiggsSignals (HS) excludes, the orange square points denote samples that only HiggsBounds (HB) excludes, and the magenta pentagon points denote samples that both HB and HS exclude.
To compare the constraints from HiggsBounds and HiggsSignals on the GNMSSM parameter space in the light Higgs scenario, the parameter samples excluded by HiggsTools in Figure 1 are color-coded in Figure 2 based on the exclusion source. The red diamond points denote samples that only HiggsSignals excluded, the orange square points denote those that only HiggsBounds excluded, and the magenta pentagon points indicate samples that both tools excluded. Figure 2 demonstrates that HiggsBounds exhibits stronger exclusion power than HiggsSignals. Almost no parameter points were only excluded by HiggsSignals (red diamonds), while points excluded by HiggsSignals are almost invariably also excluded by HiggsBounds, as indicated by the dominance of magenta points.
In the $h_{1}$ scenario, parameter points with relatively large values of $\text{Br}(h\to a_{1}a_{1}\to\tau\tau bb)$ (specifically, $\text{Br}(h\to a_{1}a_{1}\to\tau\tau bb)\geq 2.5\%$) are excluded by both HiggsSignals and HiggsBounds. Points with smaller branching ratios are only excluded by HiggsBounds. This difference occurs because HiggsBounds imposes direct constraints, whereas HiggsSignals imposes indirect constraints on non-SM-like Higgs boson searches. Specifically, if an excessively large branching ratio $\text{Br}(h\to a_{1}a_{1})$ is sufficiently large, it can affect the properties of the 125-GeV Higgs boson and implies a large $\text{Br}(h\to a_{1}a_{1}\to\tau\tau bb)$, although the reverse is not necessarily true. However, small values of $\text{Br}(h\to a_{1}a_{1}\to\tau\tau bb)$ always correspond to small $\text{Br}(h\to a_{1}a_{1})$, the latter of which has minimal impact on the properties of the 125-GeV Higgs boson.
Figure 3: Similar to the right plot in Figure 2, the label of the horizontal axis is $\text{Br}(h_{2}\to h_{1}h_{1}\to\tau\tau bb)$.
Consistent with the $h_{1}$ scenario, analogous characteristics were observed in the $h_{2}$ scenario, which simultaneously manifested distinctive features. The right panel of Figure 2 shows that some parameter points with $Br(h\to a_{1}a_{1}\to\tau\tau bb)\leq 2.5\%$ (indicated by magenta points) were also excluded by HiggsSignals, since the exotic decay channel $h\to h_{1}h_{1}$ becomes kinematically accessible in this scenario with an enhanced branching fraction. This enhancement may affect the properties of the SM-like Higgs boson and the constraints that HiggsSignals imposes. Figure 3 plots the relationship between $Br(h\to a_{1}a_{1}\to\tau\tau bb)$ and $Br(h\to h_{1}h_{1}\to\tau\tau bb)$ in the $h_{2}$ scenario. It reveals an anti-correlation between these branching ratios across the parameter space, which validates the proposed explanation.
Figure 4: Relationship between parameters $\lambda$ and $\kappa$, similar to Figure 1.
Figure 5: SM-like Higgs boson component in the $h_{1}$ scenario, similar to Figure  1.
Figure 6: $h_{2}$ component (left panel) and $h_{1}$ component (right panel) in the $h_{2}$ scenario, similar to Figure  1.
The branching ratios for the exotic decays of $h\to a_{1}a_{1}$ and $h\to h_{1}h_{1}$ depend on their respective triple Higgs couplings $C_{ha_{1}a_{1}}$ and $C_{hh_{1}h_{1}}$, as defined in Eqs. (2.18) and (2.19). These couplings are primarily governed by parameters $\lambda$ and $\kappa$ and the SM-like fraction $V_{h}^{\rm SM}$ of the $h$ boson. In the $h_{2}$ scenario, $C_{hh_{1}h_{1}}$ is also modulated by the singlet fraction $V_{h_{s}}^{\rm S}$ of the singlet-dominated state $h_{s}$. Figure 4 displays the correlation between $\lambda$ and $\kappa$ under HiggsTools constraints, which indicates that $\lambda\leq 0.1$ for most viable parameter points, which is consistent with experimental preferences for minimal doublet-singlet mixing. Figures 5 and 6 reveal the Higgs boson compositions. Both $h_{1}$ and $h_{2}$ scenarios require that the observed $h$ boson is predominantly SM-like ($V_{h}^{\rm SM}\geq 0.93$) with suppressed singlet admixture ($V_{h}^{\rm S}\leq 0.32$). For the $h_{2}$ scenario, the lighter $h_{s}$ state must exhibit high singlet purity ($V_{h_{s}}^{\rm S}\geq 0.94$) and negligible SM-like components ($V_{h_{s}}^{\rm SM}\leq 0.32$). Figure 6 further illustrates that $V_{h}^{S}\simeq-V_{h_{s}}^{\rm SM}$, confirming the theoretical derivation presented in Eq.(2.14).
Point P1 Point P2
Parameter Value Parameter Value
$\lambda$, $\kappa$, 0.015,0.3690.015,-0.369 $\lambda$, $\kappa$ 0.017,0.4550.017,-0.455
$\tan\beta$, $\delta$ $41.5,0.02$ $\tan\beta$, $\delta$ $37.6,0.08$
$v_{s}$, $\mu_{\rm tot}$, $m_{N}$ $212.8,468.9,23.6$ $v_{s}$, $\mu_{\rm tot}$, $m_{N}$ $811.7,342.0,147.7$
$m_{B}$, $m_{C}$ $208.8,144.6$ $m_{B}$$m_{C}$ $181.6,23.3$
$A_{t}$, $A_{\kappa}$ 2536.3,685.32536.3,-685.3 $A_{t}$$A_{\kappa}$ 2592.3,1152.82592.3,-1152.8
Particle Mass Spectrum Particle Mass Spectrum
$\tilde{\chi}_{1}^{0}$$\tilde{\chi}_{2}^{0}$, $\tilde{\chi}_{3}^{0}$ $23.5,478.2,483.5$ $\tilde{\chi}_{1}^{0}$$\tilde{\chi}_{2}^{0}$, $\tilde{\chi}_{3}^{0}$ $146.5,349.2,354.2$
$\tilde{\chi}_{4}^{0}$, $\tilde{\chi}_{5}^{0}$ $1008.3,2020.8$ $\tilde{\chi}_{4}^{0}$, $\tilde{\chi}_{5}^{0}$ $1008.1,2020.8$
$\tilde{\chi}_{1}^{\pm}$, $\tilde{\chi}_{2}^{\pm}$ $481.4,2021.1$ $\tilde{\chi}_{1}^{\pm}$, $\tilde{\chi}_{2}^{\pm}$ $352.2,2021.0$
$_{s}, hh$, $H$ $154.8,125.1,2414.2$ $_{s}, hh$, $H$ $188.9,125.3,2273.6$
$A_{s}$, $A_{H}$ $48.0,2414.1$ $A_{s}$, $A_{H}$ $16.6,2273.6$
Primary Decays Branching Ratios [%] Primary Decays Branching Ratios [%]
$h\to b\bar{b}$ $47.5$ $h\to b\bar{b}$ $49.4$
$h\to W^{+}W^{-}$ $22.4$ $h\to W^{+}W^{-}$ $24.0$
$h\to A_{s}A_{s}$ $16.1$ $h\to A_{s}A_{s}$ $11.9$
Rotation Matrix Element Value Rotation Matrix Element Value
$V_{h_{s}}^{S}$$V_{h_{s}}^{SM}$$V_{h}^{S}$$V_{h}^{SM}$ 0.99996, 0.0005,0.0003,0.9997-0.99996,\ 0.0005,\ -0.0003,\ -0.9997 $V_{h_{s}}^{S}$$V_{h_{s}}^{SM}$$V_{h}^{S}$$V_{h}^{SM}$ 0.99996,0.005, 0.005,0.9996-0.99996,\ -0.005,\ 0.005,\ -0.9996
$N_{11}$$N_{12}$$N_{13}$$N_{14}$$N_{15}$ 0.0003,0.0002,0.0002,0.0055,0.999980.0003,-0.0002,0.0002,-0.0055,0.99998 $N_{11}$$N_{12}$$N_{13}$$N_{14}$$N_{15}$ 0.00059, 0.0004,0.004, 0.00997,0.9999-0.00059,\ 0.0004,\ -0.004,\ 0.00997,\ -0.9999
$N_{21}$$N_{22}$$N_{23}$$N_{24}$$N_{25}$ -0.0599, 0.036, -0.71, 0.70, 0.004 $N_{21}$$N_{22}$$N_{23}$$N_{24}$$N_{25}$ 0.048, -0.033, 0.71, -0.70, -0.00996
$N_{31}$$N_{32}$$N_{33}$$N_{34}$$N_{35}$ 0.02, -0.02, -0.71, -0.71, -0.0037 $N_{31}$$N_{32}$$N_{33}$$N_{34}$$N_{35}$ -0.022, 0.023, 0.70, 0.71, 0.004
$N_{41}$$N_{42}$$N_{43}$$N_{44}$$N_{45}$ 0.998, 0.0045, -0.028, 0.056, 0 $N_{41}$$N_{42}$$N_{43}$$N_{44}$$N_{45}$ -0.998, -0.004, 0.02, -0.05, 0
$N_{51}$$N_{52}$$N_{53}$$N_{54}$$N_{55}$ 0.0019, -0.999, -0.01, 0.04, 0 $N_{51}$$N_{52}$$N_{53}$$N_{54}$$N_{55}$ 0.002, -0.999, -0.0078, 0.04, 0
Primary Annihilation Fraction [%] Primary Annihilation Fraction [%]
$\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to b\bar{b}$ $79.97$ $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to h_{s}A_{s}$ $81.2$
$\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to\tau^{+}\tau^{-}$ $19.74$ $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to A_{s}A_{s}$ $18.79$
DM Observable Value DM Observable Value
$\Omega h^{2}$ $0.000086$ $\Omega h^{2}$ $0.001593$
$\sigma^{\rm SI}_{p,n}/(10^{-49}\text{cm}^{2})$ $6.8,7.9$ $\sigma^{\rm SI}_{p,n}/(10^{-48}\text{cm}^{2})$ $4.1,4.8$
$\sigma^{\rm SD}_{p,n}/(10^{-50}\text{cm}^{2})$ $2.5,1.9$ $\sigma^{\rm SD}_{p,n}/(10^{-48}\text{cm}^{2})$ $3.6,2.8$
Table 2: Benchmark points P1 and P2 compatible with all experimental constraints. Both points correspond to Singlino-dominated DM in the $h_{1}$ scenario. All mass parameters are given in GeV.
Point P3 Point P4
Parameter Value Parameter Value
$\lambda$, $\kappa$, $0.16,0.117$ $\lambda$, $\kappa$ 0.026,0.0160.026,-0.016
$\tan\beta$, $\delta$ 20.4,0.0220.4,-0.02 $\tan\beta$, $\delta$ $23.7,0.108$
$v_{s}$, $\mu_{\rm tot}$, $m_{N}$ $739.4,269.5,776.5$ $v_{s}$, $\mu_{\rm tot}$, $m_{N}$ 872.6,405.6,655.4872.6,405.6,-655.4
$m_{B}$, $m_{C}$ $282.3,240.0$ $m_{B}$$m_{C}$ $51.2,73.5$
$A_{t}$, $A_{\kappa}$ 2694.0,1683.32694.0,-1683.3 $A_{t}$$A_{\kappa}$ $2720.2,1859.9$
Particle Mass Spectrum Particle Mass Spectrum
$\tilde{\chi}_{1}^{0}$$\tilde{\chi}_{2}^{0}$, $\tilde{\chi}_{3}^{0}$ $273.8,279.6,777.0$ $\tilde{\chi}_{1}^{0}$$\tilde{\chi}_{2}^{0}$, $\tilde{\chi}_{3}^{0}$ $413.0,418.1,655.4$
$\tilde{\chi}_{4}^{0}$, $\tilde{\chi}_{5}^{0}$ $1008.1,2020.8$ $\tilde{\chi}_{4}^{0}$, $\tilde{\chi}_{5}^{0}$ $1008.2,2020.7$
$\tilde{\chi}_{1}^{\pm}$, $\tilde{\chi}_{2}^{\pm}$ $277.3,2021.0$ $\tilde{\chi}_{1}^{\pm}$, $\tilde{\chi}_{2}^{\pm}$ $416.1,2021.0$
$_{s}, hh$, $H$ $169.1,125.3,2104.7$ $_{s}, hh$, $H$ $21.77,125.4,2177.9$
$A_{s}$, $A_{H}$ $61.2,2103.9$ $A_{s}$, $A_{H}$ $56.8,2177.9$
Primary Decays Branching Ratios [%] Primary Decays Branching Ratios [%]
$h\to b\bar{b}$ $52.1$ $h\to b\bar{b}$ $48.4$
$h\to W^{+}W^{-}$ $24.0$ $h\to W^{+}W^{-}$ $23.7$
$h\to A_{s}A_{s}/h_{s}h_{s}$ $8.8/0$ $h\to A_{s}A_{s}/h_{s}h_{s}$ $3.6/9.96$
Rotation Matrix Element Value Rotation Matrix Element Value
$V_{h_{s}}^{S}$$V_{h_{s}}^{SM}$$V_{h}^{S}$$V_{h}^{SM}$ 0.999, 0.03,0.03,0.998-0.999,\ 0.03,\ -0.03,\ -0.998 $V_{h_{s}}^{S}$$V_{h_{s}}^{SM}$$V_{h}^{S}$$V_{h}^{SM}$ 0.9997,0.021, 0.021,0.99890.9997,\ -0.021,\ 0.021,\-0.9989
$N_{11}$$N_{12}$$N_{13}$$N_{14}$$N_{15}$ 0.044,0.033,0.71,0.70,0.0370.044,-0.033,0.71,-0.70,0.037 $N_{11}$$N_{12}$$N_{13}$$N_{14}$$N_{15}$ 0.054,0.035, 0.71,0.70,0.00290.054,\ -0.035,\ 0.71,\ -0.70,\ -0.0029
$N_{21}$$N_{22}$$N_{23}$$N_{24}$$N_{25}$ 0.023, 0.023, 0.70, 0.71, 0.019-0.023,\ 0.023,\ 0.70,\ 0.71,\ 0.019 $N_{21}$$N_{22}$$N_{23}$$N_{24}$$N_{25}$ 0.021,0.022,$0.71,$ 0.0140.021,\ -0.022,\ -0.71,\ -0.71,\ 0.014
$N_{31}$$N_{32}$$N_{33}$$N_{34}$$N_{35}$ 0.0001, 0.0009,0.04, 0.01, 0.999-0.0001,\ 0.0009,\ -0.04,\ 0.01,\ 0.999 $N_{31}$$N_{32}$$N_{33}$$N_{34}$$N_{35}$ 0.0001, 0.0002, 0.012, 0.0078, 0.999899-0.0001,\ 0.0002,\ 0.012,\ 0.0078,\ 0.999899
$N_{41}$$N_{42}$$N_{43}$$N_{44}$$N_{45}$ 0.9987,0.004, 0.015,0.047, 0.001-0.9987,\ -0.004,\ 0.015,\ -0.047,\ 0.001 $N_{41}$$N_{42}$$N_{43}$$N_{44}$$N_{45}$ 0.998,0.004, 0.02,0.05, 0-0.998,\ -0.004,\ 0.02,\ -0.05,\ 0
$N_{51}$$N_{52}$$N_{53}$$N_{54}$$N_{55}$ 0.002, 0.999, 0.007,0.039,0.0001-0.002,\ 0.999,\ 0.007,\ -0.039,\ -0.0001 $N_{51}$$N_{52}$$N_{53}$$N_{54}$$N_{55}$ 0.002,0.999,0.0098, 0.04, 00.002,\ -0.999,\ -0.0098,\ 0.04,\ 0
Primary Annihilation Fraction [%] Primary Annihilation Fraction [%]
$\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{\pm}\to d\bar{u}/\bar{d}u$ $10.58$ $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{\pm}\to d\bar{u}/\bar{d}u$ $10.05$
$\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{\pm}\to s\bar{c}/\bar{s}c$ $10.58$ $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{\pm}\to s\bar{c}/\bar{s}c$ $10.05$
$\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{\pm}\to b\bar{t}/t\bar{b}$ $7.8$ $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{\pm}\to b\bar{t}/t\bar{b}$ $6.66$
$\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to W^{+}W^{-}$ $6.7$ $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to W^{+}W^{-}$ $4.84$
$\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to ZZ$ $5.21$ $\tilde{\chi}_{2}^{0}\tilde{\chi}_{1}^{\pm}\to d\bar{u}/\bar{d}u$ $4.41$
$\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{\pm}\to e\bar{\nu_{e}}/\bar{\nu_{e}}e$ $3.53$ $\tilde{\chi}_{2}^{0}\tilde{\chi}_{1}^{\pm}\to s\bar{c}/\bar{s}c$ $4.41$
$\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{\pm}\to\mu\bar{\nu_{\mu}}/\bar{\nu_{\mu}}\mu$ $3.53$ $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to ZZ$ $3.81$
$\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{\pm}\to\tau\bar{\nu_{\tau}}/\bar{\nu_{\tau}}\mu$ $3.50$ $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{\pm}\to e\bar{\nu_{e}}/\bar{\nu_{e}}e$ $3.35$
$\tilde{\chi}_{2}^{0}\tilde{\chi}_{1}^{\pm}\to d\bar{u}/\bar{d}u$ $3.17$ $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{\pm}\to\mu\bar{\nu_{\mu}}/\bar{\nu_{\mu}}\mu$ $3.35$
$\tilde{\chi}_{2}^{0}\tilde{\chi}_{1}^{\pm}\to s\bar{c}/\bar{s}c$ $3.17$ $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{\pm}\to\tau\bar{\nu_{\tau}}/\bar{\nu_{\tau}}\tau$ $3.27$
DM Observable Value DM Observable Value
$\Omega h^{2}$ $0.0091$ $\Omega h^{2}$ $0.02$
$\sigma^{\rm SI}_{p,n}/(10^{-48}\text{cm}^{2})$ $6.32,5.47$ $\sigma^{\rm SI}_{p,n}/(10^{-48}\text{cm}^{2})$ $1.78,4.85$
$\sigma^{\rm SD}_{p,n}/(10^{-43}\text{cm}^{2})$ $1.64,1.26$ $\sigma^{\rm SD}_{p,n}/(10^{-43}\text{cm}^{2})$ $4.08,3.13$
Table 3: Benchmark points P3 and P4 that satisfy all experimental constraints. Points P3 and P4 represent Higgsino-dominated DM in the $h_{1}$ and $h_{2}$ scenarios, respectively. All mass parameters are given in GeV.
3.3 Dark matter physics
To further illustrate the features of the GNMSSM parameter space in the light Higgs scenario, DM phenomenology was investigated. The package MicrOMEGAs  [81, 82, 83, 84, 85, 86, 87] was used to calculate the DM relic density and the spin-dependent (SD) and spin-independent (SI) DM-nucleon cross sections. It was assumed that a large DM population existed in the early universe and froze out to match the current Planck observation, $\Omega_{\rm DM}h^{2}=0.12\pm 0.01$ [88]. In this study, the relic density of DM (the lightest neutralino $\tilde{\chi}_{1}^{0}$) was required to be below the central value of 0.12. The SI and SD cross sections were scaled by a factor of $\Omega h^{2}/0.12$. It was found that in the $h_{1}$ scenario, the DM can be either Singlino- or Higgsino-dominated, whereas in the $h_{2}$ scenario, it is Higgsino-dominated.
Four benchmark points were selected, as listed in Tables 2 and 3, based on their dominant annihilation processes. These points also satisfy the 2024 LZ  [30] direct detection constraints. Points P1 and P2 correspond to Singlino-dominated DM in the $h_{1}$ scenario. After applying the LZ-2024 constraints, the values of $\lambda$ are generally small. For Singlino-dominated DM, $m_{\tilde{\chi}_{1}^{0}}\simeq m_{N}$, consistent with Eq.(2.17). For benchmark point P1, mostly annihilation channels of DM are $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to b\bar{b}/\tau^{+}\tau^{-}$. In contrast, for point P2, since $2m_{\tilde{\chi}_{1}^{0}}>m_{h_{s}}+m_{A_{s}}$, the channel $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to h_{s}A_{s}$ is kinematically allowed and mostly annihilation channel of DM. This process occurs via s-channel exchange of CP-odd Higgs bosons and t-channel exchange of neutralinos. Additionally, the channel $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to A_{s}A_{s}$ is also open, proceeding through s-channel CP-even Higgs exchange and t-channel neutralino exchange.
Points P3 and P4 correspond to Higgsino-dominated DM in the $h_{1}$ and $h_{2}$ scenarios, respectively. For these benchmark points, mostly annihilation channels of DM are achieved through co-annihilation processes $\tilde{\chi}_{i}^{0}\tilde{\chi}_{1}^{\pm}\to XY$ ($i=1,2$), where X and Y represent quarks or leptons in SM, as well as through the annihilation channels $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to W^{+}W^{-}/ZZ$. In the Higgsino-dominated DM scenario, the masses of the lightest neutralino, the next-to-lightest neutralino, and the lightest chargino are nearly degenerate. Given the complexity of DM physics and its close relation to Higgs phenomenology, a detailed study of DM properties, particularly in the light Higgs scenario, will be the focus of future work.
4 Conclusion
This study systematically investigated the exotic decay channels of the standard-model-like Higgs boson into pairs of the lightest CP-odd or CP-even Higgs bosons within the framework of GNMSSM. First, we performed a comprehensive scan of the GNMSSM parameter space using the MultiNest algorithm, subject to experimental constraints from HiggsSignals-2.6.2 and HiggsBounds-5.10.2. To assess the implications of recent Higgs precision measurements for GNMSSM scenarios featuring light Higgs states (considering both cases where $h_{1}=h_{\text{SM}}$ and $h_{2}=h_{\text{SM}}$), we separately examined the exclusion limits provided by HiggsSignals and the HiggsBounds implementation in HiggsTools. Within the phenomenologically allowed parameter region, we further conducted a preliminary analysis of dark matter properties. The main findings were summarized as follows:
  • In both the $h_{1}$ and $h_{2}$ scenarios, HiggsBounds imposes the most stringent constraints on the theoretical parameter space. In contrast, HiggsSignals and individual ATLAS searches exhibit comparatively weaker constraining power; moreover, the vast majority of parameter points excluded by these analyses are also disfavored by HiggsBounds . This is because HiggsBounds provides direct constraints from dedicated searches for non-Standard Model Higgs bosons, whereas HiggsSignals derives indirect constraints based on global compatibility with experimental data.
  • It is worth noting that the analysis of the $h_{2}$ scenario reveals unique phenomenological features. Specifically, HiggsSignals can exclude certain parameter points with relatively low values for certain exotic decay branches, especially when $Br(h\to a_{1}a_{1}\to\tau\tau bb)\leq 2.5\%$. This exclusion effect stems from the kinematic accessibility of the $h_{2}\to h_{1}h_{1}$ decay process, which has a significantly enhanced branching ratio. Such enhanced decays will modify the properties of the standard model-like Higgs boson, thereby increasing the sensitivity of HiggsSignals to indirect constraints in regions where they were previously suppressed due to kinematic factors.
  • Under the combined constraints of HiggsTools, both $h_{1}$ and $h_{2}$ scenarios require that the observed $h$ boson is predominantly SM-like ($V_{h}^{\rm SM}\geq 0.93$) with minimal singlet admixture ($V_{h}^{\rm S}\leq 0.32$). In the $h_{2}$ scenario, the lighter $h_{s}$ state must exhibit high singlet purity ($V_{h_{s}}^{\rm S}\geq 0.94$) and negligible SM-like components ($V_{h_{s}}^{\rm SM}\leq 0.32$).
  • In the $h_{1}$ scenario, DM can be either Singlino- or Higgsino-dominated, whereas in the $h_{2}$ scenario, it is Higgsino-dominated. For Higgsino-dominated DM, mostly annihilation channels of DM are achieved through co-annihilation with chargino; For Singlino-dominated DM, mostly annihilation channels of DM may be $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to h_{s}A_{s}$.
Despite constraints from current Higgs data on light Higgs scenarios, phenomenologically viable regions in the parameter space persist. These regions merit further investigation for two principal reasons. First, dark matter physics exhibits considerable complexity and is deeply intertwined with Higgs phenomenology. Second, the corresponding mass spectrum can induce a strong first-order electroweak phase transition (SFOEWPT) in the early universe [89, 90]. This mechanism satisfies the necessary conditions for electroweak baryogenesis (EWBG), providing a plausible explanation for the observed baryon asymmetry of the universe, and simultaneously produces detectable gravitational wave (GW) signals during the phase transition. These aspects will form the core focus of our future work.
Acknowledgement
We sincerely thank Prof. Junjie Cao for helpful discussions. We thank LetPub [91] for its linguistic assistance during the preparation of this manuscript. This work was supported by the National Natural Science Foundation of China (Grant No. 12505122).